Find the equation of the tangent and normal lines to the curve y3-xy+x2=1 at (1,0)

1
Expert's answer
2022-05-31T09:12:50-0400
"y^3-xy+x^2=1"

Differentiate both sides with respect to "x"


"3y^2y'-y-xy'+2x=0"

"y'=\dfrac{y-2x}{3y^2-x}"

"slope=m=\dfrac{0-2(1)}{3(0)^2-1}=2"

The tangent line in point-slope form is


"y-0=2(x-1)"

The tangent line in slope-intercept form is


"y=2x-2"


For the normal line


"slope_2=m_2=-\dfrac{1}{m}=-\dfrac{1}{2}"

The normal line in point-slope form is


"y-0=-\dfrac{1}{2}(x-1)"

The normal line in slope-intercept form is


"y=-\dfrac{1}{2}x+\dfrac{1}{2}"